OpenStudy (anonymous):

slope-intercept form (7,9):perpendicular to x=7

OpenStudy (anonymous):

so perpendicular means the slope must be the negative reciprocal, so if the slope of x=7 is 1, then the slope of the perpendicular line must be -1/1, or just -1.

OpenStudy (anonymous):

so we have y=mx+b y=-1x+b

OpenStudy (anonymous):

plug in the point (7,9) into the function for x & y and solve for b

OpenStudy (anonymous):

then place into the form y=mx+b

OpenStudy (anonymous):

still having trouble

OpenStudy (anonymous):

so how far can you get on your own?

OpenStudy (anonymous):

i got it to the plug in the numbers now i dont know to to put in y=mx+b

OpenStudy (anonymous):

so what do you find for b?

OpenStudy (anonymous):

b=16 right

OpenStudy (anonymous):

yup so you have m=-1, and b=16, and x and y are just variables

OpenStudy (anonymous):

so y=-x+16

OpenStudy (anonymous):

oh ok i thought i had to have number also than you

OpenStudy (anonymous):

nope!